# Beyond Score Changes: Adversarial Attack on No-Reference Image Quality Assessment from Two Perspectives

### v1

@@ 3.4. Targeted Samples Generation
(12) $\min_{I^{\prime}_{j}}\left(f(I^{\prime}_{j})-s^{*}_{j}\right)^{2},~{}~{}~{}~{}\text{s.t. }\left\|I^{\prime}_{j}-I_{j}\right\|_{\infty}\leqslant\epsilon.$
@@
To efficiently generate $I^{\prime}_{j}$ satisfying Eq. (12), we draw inspiration from a classical attack method in the classification task — the fast gradient sign method (FGSM) (Goodfellow et al., 2015). We adopt an iterative approach to update the adversarial sample, incorporating the clamp function to ensure that $\|I^{\prime}_{j}-I_{j}\|\infty\leqslant\epsilon$.
@@
(13) $I^{\prime}_{j}\leftarrow I^{\prime}_{j}-\alpha\cdot\text{sign}\left(\nabla_{I^{\prime}_{j}}\left(f(I^{\prime}_{j})-s^{*}_{j}\right)^{2}\right)$
@@ 4. Experiments
In this section, we present the experimental results conducted on the CLIVE (Ghadiyaram and Bovik, 2016) dataset, which collected images distorted in the wild and is challenging for the NR-IQA task. We begin by comparing the proposed SMA method with four state-of-the-art (SOTA) attack methods on four NR-IQA models in Sec. 4.1. Then we analyze the insight provided by the optimization strategy used in Stage One in Sec. 4.2. We discuss the error between the target scores and predicted scores in Stage Two in Sec. 4.3. Ablation studies that explore hyperparameters in this work are presented in Sec. 4.4.
@@ 4.1. Comparison with SOTA Attacks
We compare our method with four NR-IQA attack methods: the perceptual attack (P-attack) (Zhang et al., 2022), OUAP (Shumitskaya et al., 2024), the attack method proposed by Korhonen and You (K-attack) (2022), and UAP (Shumitskaya et al., 2022). Among them, K-attack and UAP generate adversarial examples on a substitute model and then transfer these samples to the target NR-IQA model. We select four widely-used NR-IQA models as the target model: DBCNN (Zhang et al., 2020), HyperIQA (Su et al., 2020), MANIQA (Yang et al., 2022), and LIQE (Zhang et al., 2023b), whose architectures are based on ResNet (He et al., 2016), VGG (Simonyan and Zisserman, 2015), transformer ViT (Dosovitskiy et al., 2021) and CLIP (Radford et al., 2021), respectively. For all models, $80\%$ of the images are randomly selected from the CLIVE dataset for training NR-IQA models, with the remaining images (a total of $232$ images) for testing. During the attack phase, only the test images are used. According to the original settings of these models, images with full size $500\times 500$ are the inputs of DBCNN, and images are center-cropped to the size of $224\times 224$ before being input to other three NR-IQA models. Additionally, we add a clipping function into the final layer of all NR-IQA models to ensure that predicted scores fall within the range of $[0,100]$.
@@
(14) $R=\frac{1}{N}\sum_{j=1}^{N}\log\left(\frac{\max\{\beta_{1}-f(I_{j}),f(I_{j})-\beta_{2}\}}{|f(I_{j})-f(I_{j}^{\prime})|}\right),$
@@
(15) $\Delta\text{Rank}=\frac{1}{N}\sum_{j=1}^{N}\left|r_{j}(f(\mathcal{I^{\prime}}))-r_{j}(f(\mathcal{I}))\right|,$
@@
Meanwhile, we consider the imperceptibility of adversarial perturbations, ensuring that the MOS of adversarial examples is the same as that of clean images. We employ the commonly-used metrics, LPIPS (Zhang et al., 2018), to quantify the visual similarity between adversarial examples and clean images. Smaller LPIPS values indicate more imperceptible perturbations.
@@
In our experiments, we use $r^{1}(\cdot)$ to approximate the ranking function. $\lambda_{\text{Var}},\lambda_{\text{MSE}}$ in Eq. (11) are chosen as $0.0001$ for DBCNN, HyperIQA and LIQE. For MANIQA, $\lambda_{\text{Var}},\lambda_{\text{MSE}}$ are set to $0.0002$ and $0.00001$, respectively. The objective in Eq. (11) is optimized $100,000$ epochs with Adam optimizer (Kingma and Ba, 2015). The $\epsilon$ in Eq. (12) is set to $0.005$. The iteration number $m$ and step size $\alpha$ are set to $10,0.005$, respectively. The comparative results are shown in Table 1.
@@ 4.4. Ablation Study
There are three main hyperparameters in SMA: $\beta$ in the approximation function $r^{\beta}(\cdot)$, the multiplier $\lambda_{\text{Var}}$ and $\lambda_{\text{MSE}}$ in Eq. (11). This section presents ablation studies on them with the target model being HyperIQA.
@@
The weight $\lambda_{\text{MSE}}$ in Eq. (11). A larger $\lambda_{\text{MSE}}$ implies a more strict constraint on the error between target scores and original scores, thereby enhances the error between original scores and predicted scores of adversarial examples generated in Stage Two. Fig. 8 (right) illustrates the impact of $\lambda_{\text{MSE}}$ when $\lambda_{\text{Var}}=0$. As $\lambda_{\text{MSE}}$ increases, the RMSE between $\boldsymbol{s}$ and $f(\mathcal{I^{\prime}})$ also increases, confirming the effectiveness of $\text{MSE}(\cdot,\cdot)$ in boosting the error between predicted scores before and after the attack. Meanwhile, the increase of $\lambda_{\text{MSE}}$ will result in a decrease of SROCC to some extent.
@@
Figure 8. Ablation studies of the weight $\lambda_{\text{Var}}$ (left) and $\lambda_{\text{MSE}}$ (right) in Eq. (11).

### v2

@@ 3.4. Targeted Samples Generation
$\min_{I^{\prime}_{j}}\left(f(I^{\prime}_{j})-s^{*}_{j}\right)^{2},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{s.t. }\left\|I^{\prime}_{j}-I_{j}\right\|_{\infty}\leqslant\epsilon.$ (12)
@@
To efficiently generate $I^{\prime}_{j}$ satisfying Eq. (12), we draw inspiration from a classical attack method in the classification task — the fast gradient sign method (FGSM) Goodfellow et al. (2015). We adopt an iterative approach to update the adversarial sample, incorporating the clamp function to ensure that $\|I^{\prime}_{j}-I_{j}\|\infty\leqslant\epsilon$.
@@
$I^{\prime}_{j}\leftarrow I^{\prime}_{j}-\alpha\cdot\text{sign}\left(\nabla_{I^{\prime}_{j}}\left(f(I^{\prime}_{j})-s^{*}_{j}\right)^{2}\right),$ (13)
@@ 4. Experiments
In this section, we present the experimental results conducted on the CLIVE Ghadiyaram and Bovik (2016) dataset, which collected images distorted in the wild and is challenging for the NR-IQA task. We begin by comparing the proposed SMA method with four state-of-the-art (SOTA) attack methods on four NR-IQA models in Sec. 4.1. Then we analyze the insight provided by the optimization strategy used in Stage One in Sec. 4.2. We discuss the error between the target scores and predicted scores in Stage Two in Sec. 4.3. Ablation studies that explore hyperparameters in this work are presented in Sec. 4.4.
@@ 4.1. Comparison with SOTA Attacks
We compare our method with four NR-IQA attack methods: the perceptual attack (P-attack) Zhang et al. (2022), OUAP Shumitskaya et al. (2024), the attack method proposed by Korhonen and You (K-attack) Korhonen and You (2022), and UAP Shumitskaya et al. (2022). Among them, K-attack and UAP generate adversarial examples on a substitute model and then transfer these samples to the target NR-IQA model. We select four widely-used NR-IQA models as the target model: DBCNN Zhang et al. (2020), HyperIQA Su et al. (2020), MANIQA Yang et al. (2022), and LIQE Zhang et al. (2023b), whose architectures are based on ResNet He et al. (2016), VGG Simonyan and Zisserman (2015), transformer ViT Dosovitskiy et al. (2021) and CLIP Radford et al. (2021), respectively. For all models, $80\%$ of the images are randomly selected from the CLIVE dataset for training NR-IQA models, with the remaining images (a total of $232$ images) for testing. During the attack phase, only the test images are used. According to the original settings of these models, images with full size $500\times 500$ are the inputs of DBCNN, and images are center-cropped to the size of $224\times 224$ before being input to other three NR-IQA models. Additionally, we add a clipping function into the final layer of all NR-IQA models to ensure that predicted scores fall within the range of $[0,100]$.
@@
$R=\frac{1}{N}\sum_{j=1}^{N}\log\left(\frac{\max\{\beta_{1}-f(I_{j}),f(I_{j})-\beta_{2}\}}{|f(I_{j})-f(I_{j}^{\prime})|}\right),$ (14)
@@
$\Delta\text{Rank}=\frac{1}{N}\sum_{j=1}^{N}\left|r_{j}(f(\mathcal{I^{\prime}}))-r_{j}(f(\mathcal{I}))\right|,$ (15)
@@
Meanwhile, we consider the imperceptibility of adversarial perturbations, ensuring that the MOS of adversarial examples is the same as that of clean images. We employ the commonly-used metrics, LPIPS Zhang et al. (2018), to quantify the visual similarity between adversarial examples and clean images. Smaller LPIPS values indicate more imperceptible perturbations.
@@
In our experiments, we use $r^{1}(\cdot)$ to approximate the ranking function. $\lambda_{\text{Var}},\lambda_{\text{MSE}}$ in Eq. (11) are chosen as $0.0001$ for DBCNN, HyperIQA and LIQE. For MANIQA, $\lambda_{\text{Var}},\lambda_{\text{MSE}}$ are set to $0.0002$ and $0.00001$, respectively. The objective in Eq. (11) is optimized $100,000$ epochs with Adam optimizer Kingma and Ba (2015). The $\epsilon$ in Eq. (12) is set to $0.005$. The iteration number $m$ and step size $\alpha$ are set to $10,0.005$, respectively. The comparative results are shown in Table 1.
@@ 4.4. Ablation Study
There are three main hyperparameters in SMA: $\beta$ in the approximation function $r^{\beta}(\cdot)$, the multiplier $\lambda_{\text{Var}}$, and $\lambda_{\text{MSE}}$ in Eq. (11). We present ablation studies on them with the target model HyperIQA.
@@
Figure 8. Ablation studies of the multiplier $\lambda_{\text{Var}}$ and $\lambda_{\text{MSE}}$.
The weight $\lambda_{\text{MSE}}$ in Eq. (11). A larger $\lambda_{\text{MSE}}$ implies a more strict constraint on the error between target scores and original scores, thereby enhancing the error between original scores and predicted scores of adversarial examples generated in Stage Two. Fig. 8 (right) illustrates the impact of $\lambda_{\text{MSE}}$ when $\lambda_{\text{Var}}=0$. As $\lambda_{\text{MSE}}$ increases, the RMSE between $\boldsymbol{s}$ and $f(\mathcal{I^{\prime}})$ also increases, confirming the effectiveness of $\text{MSE}(\cdot,\cdot)$ in boosting the error between predicted scores before and after the attack. Meanwhile, the increase of $\lambda_{\text{MSE}}$ will decrease SROCC to some extent.